# Together We Go Further: LLMs and IDE Static Analysis for Extract Method Refactoring

### v1

@@ 3. Technique
Figure 2. The workflow of generating refactoring suggestions and then applying them with EM-Assist.
In this section, we present the workflow that our novel approach and tool, EM-Assist, uses to automatically suggest and perform Extract Method refactoring for long methods in Java and Kotlin codebases. Figure 2 shows the architecture and the steps performed by EM-Assist. First, EM-Assist invokes the LLM by prompting it to generate Extract Method suggestions for the selected long methods ( 1 in Figure 2). We further elaborate on this in Section 3.1. Following this, our algorithm filters out hallucinations and retains applicable refactoring options that are free of errors ( 2 in Figure 2) and are useful ( 3 in Figure 2). We discuss these in detail in Section 3.2 and Section 3.3. These steps play a key role in selecting realistic suggestions that align with the way developers extract code, while simultaneously mitigating the risk of introducing bugs or non-compilable code. Next, our algorithm further enhances ( 4 in Figure 2) the quality of the suggestions by expanding or shrinking the code fragment in the suggestion based on program slicing rules presented in Section 3.4. These rules ensure that refactorings better align with the developer’s intentions. As there could be several alternatives (which can overwhelm the user) for how to split a long method, EM-Assist ranks the candidates ( 5 in Figure 2) (see Section 3.5), before it presents a high-quality subset to the user. The GUI displays the suggestions and shows a preview of the signature of the extracted method, along with the code fragment that would be extracted. Based on the user’s decision, EM-Assist encapsulates the suggestion in a refactoring command and invokes IntelliJ IDEA to perform the refactoring correctly. You can find EM-Assist, instructions for its installment, and example of its working in the supplementary material (Anonymous, 2023).
Definition 3.0 ().
(Long method) denoted as $l_{i}(l_{s},l_{e})$, in a codebase is quantified as $L(l_{i})$, where $L(l_{i})$ represents the length of the method. This length is determined as the difference between the starting line number of the method in the file ($l_{s}$) and the ending line number of the method ($l_{e}$), accounting for spaces and comments while excluding Javadoc comments. A method $l_{i}$ is classified as “long” if it exceeds a predefined threshold, such that $L(l_{i})>TV$, with $TV$ signifying the specified threshold value. We determine this threshold value empirically, and we use it to define the minimum length at which a method is considered “long” within the codebase.
Definition 3.0 ().
@@
(Extract Method) denoted as $e_{i}(n_{i},(s_{i},f_{i}))$, where $n_{i}$ represents the method name, $(s_{i},f_{i})$ signifies a code fragment within a long method $l_{i}$. It is characterized by the starting line number, defined as $s_{i}>l_{s}$, and the ending line number, defined as $e_{i}<l_{e}$. These line numbers are calculated with respect to the original source code file in which the host method $l_{i}$ is located.
Definition 3.0 ().
@@
(Extract Method Suggestions ($\mathcal{S}$)) is a set of extract method suggestions generated by an LLM for a long method $l_{m}$, formally, $\mathcal{S}=\{e_{i}(n_{i},(s_{i},f_{i})),e_{j}(n_{j},(s_{j},f_{j})),\ldots\}$
Definition 3.0 ().
@@
(Not Useful Extract Methods ($\mathcal{NU}$)) is a subset of $\mathcal{V}$ that does not satisfy the usefulness conditions ($U(e_{i})$), ensuring that the suggestions are neither too large, encompassing almost the entire method body, nor too small, essentially one-liners. Formally, $\mathcal{NU}_{i}=\{e_{i}\in\mathcal{V}\mid\neg U(e_{i})\}$.
Definition 3.0 ().
(Applicable Extract Methods ($\mathcal{A}$)) constitute a set of Extract Method suggestions within a long method $l_{i}$ that have been filtered to exclude both invalid and not useful suggestions. Formally, $\mathcal{A}_{i}=\{e_{i}\in\mathcal{S}\mid e_{i}\notin(\mathcal{I}_{i}\cup\mathcal{NU}_{i})\}.$
3.1. Generating Extract Method Suggestions
@@
We utilize LLMs to generate Extract Method recommendations for a given input long method. In this section, we delve into the preparation of LLM for Extract Method refactoring suggestions, as well as the parameters of the LLM that require tuning.
3.1.1. Prompt engineering
@@
LLMs are versatile models capable of various applications, but they require specific preparation when used for a particular task (Sidong and Chunyang, 2024). To facilitate this, we employ in-context learning (Wei et al., 2022), where we provide all the instructions needed to solve the task right in the model’s input, including task definition and relevant context information. To further enrich the prompt, we employ few-shot learning (Brown et al., 2020; Radford et al., 2019) by incorporating a set of 2 examples into the prompt, following best practices as discussed by Gao et al. (2023). The prompt includes: (i) A succinct overview of Extract Method refactoring, (ii) The definition of a long method, (iii) A JavaDoc string when available, (iv) 2 examples of refactoring for one long and one short method, and (v) Precise instructions for the desired output format. An example prompt is accessible on our companion website (Anonymous, 2023).
3.1.2. Generating an extensive array of suggestions.
@@
The output of an LLM depends on two factors: (i) the internal variable “Temperature” ($T$), and (ii) the number of iterations ($I$(Brown et al., 2020; Liu et al., 2020; Radford et al., 2019), which indicates how many times the same prompts are inputted. Temperature serves as a regulator for the model’s output randomness. Higher values, such as 0.9 or 1.2, produce more diverse and unpredictable outputs, whereas values closer to 0 produce focused and less diverse results. We determine the optimal value empirically (see Section 4.3).
@@
Even when identical prompts are presented multiple times, the output can exhibit variations due to a couple of factors: (i) LLMs, such as GPT-3.5 and GPT-4, employ a combination of learned patterns and random sampling in generating responses, introducing slight deviations in their answers each time due to inherent randomness; and (ii) LLMs possess the capability to explore various potential solutions and refine their responses iteratively through interactions and feedback (Brown et al., 2020; Radford et al., 2019), continually enhancing their outputs in subsequent iterations. Therefore, the quality and quantity of predictions made by an LLM are influenced by the temperature setting and the frequency of prompting the same prompt. In Section 4.3, we conduct a sensitivity analysis to determine how the performance of EM-Assist varies with these settings. This analysis enables EM-Assist to utilize the optimal settings for generating refactoring suggestions.
3.2. Removing Invalid Extract Method Suggestions
Not all the suggestions generated by LLMs can be directly applied to codebases. Some suggestions may result in non-compilable code (e.g., ( 3 in our motivating example from Figure 1), which are hallucinations that developers are unlikely to accept (Patil et al., 2023).
@@ 3.2. Removing Invalid Extract Method Suggestions
Scope Analysis: As the first check, EM-Assist analyzes the suggestion to ensure that it can be extracted without breaking the code’s scope. The code should be self-contained, with all necessary variables and objects accessible within the same scope or passed as parameters. For example, suggestion 3 in our motivating example (Figure 1) fails the scope analysis, and EM-Assist expands the code fragment of the suggestion so that both the start and end lines have the same scoping level: the new suggestion’s code fragment starts from line 160, while the end line stays the same.
@@
To remove suggestions that might lead to non-compilable code if applied as Extract Method, EM-Assist uses the rules below:
@@
    Return Values: If the code fragment to be extracted produces a result or modifies the state of objects, EM-Assist validates that the return value and side effects are appropriately handled. For example, it checks if the return value is used subsequently in the host method or if any thrown exceptions are caught.
@@
  3. (iii)
    Control Flow: EM-Assist reasons about the control flow within the code fragment in the suggestion to ensure that it can be extracted as a separate method without causing logical errors. For example, it checks for any early returns or breaks that might affect the control flow and would render the extracted method to have a different behavior than the original code.
EM-Assist checks all these conditions through the static analysis infrastructure in the IntelliJ Platform [^1] The IntelliJ Platform: https://www.jetbrains.com/opensource/idea/. (the open-source platform IntelliJ IDEA and other JetBrains IDEs are built upon) that verifies refactoring preconditions. As LLMs are not aware of the semantics of Extract Method refactoring, but only about the code tokens that are associated together in its training data, this step is crucial for discarding a large number of invalid suggestions (45.7% of all suggestions in our formative study).
3.3. Removing Refactoring Suggestions That Are Not Useful
Once we retain only valid refactoring suggestions, the next step is to filter out those suggestions that are not useful and retain the ones that are applicable and likely to be performed by a developer. To identify suggestions that are not useful, we employ two rules: (i) Exclude Extract Method suggestions that encompass 86% or more of the statements present in the long method. (ii) Following practices adopted by other researchers (Tsantalis and Chatzigeorgiou, 2011; Fernandes et al., 2022; Haas and Hummel, 2015; Yang et al., 2009b), we exclude Extract Method suggestions that are one-line code extractions.
Extracting an entire method or a one-liner into another method is, in general, not useful for the purpose of improving the existing code. This might be useful if the developers plan to introduce new code statements and implement new features into the host method or the extracted method that expand the original functionality. Notice that we designed EM-Assist for code renovation and not for the feature expansion use case.
3.4. Enhancing Refactoring Suggestions
After identifying applicable suggestions devoid of hallucinations, the subsequent step involves refining and enhancing these selections. During this phase, we apply the following enhancements to further improve the quality of the valid suggestions.
@@ 3.4. Enhancing Refactoring Suggestions
    Program slicing: Taking inspiration from tools that rely solely on static analysis, such as those utilizing program slicing (Maruyama, 2001; Tsantalis and Chatzigeorgiou, 2011; Charalampidou et al., 2017; Abadi et al., 2009; Meananeatra et al., 2018), we leverage program slicing to augment refactoring suggestions provided by LLMs. We use rules that better align with developer preferences to avoid creating small methods with several arguments.
    Let $EF_{s}$ represent the first statement of the code fragment to be extracted into a new method. Let $EF_{s-1}$ represent a variable declaration statement that initializes a variable $v$ that is used in the code fragment of the suggested extract method. $v$ is called a live variable, determined through a def-use chain. In such cases, we include $EF_{s-1}$ in the extracted method, and we adjust the starting line of the suggestion to match the starting line of $EF_{s-1}$. This avoids having to pass $v$ as a parameter to the extracted method. In the future, we plan to experiment with other such heuristics based on the feedback we receive from users of EM-Assist.
@@
    Control statements: If the code fragment in the extracted method starts with a control statement (i.e., $EF_{s}$ is the check block of an if statement), we shrink the code fragment to start at $EF_{s+1}$ and contain only the block of the if. This leaves the check condition intact in the host method so that the developer can see under which conditions the would-be extracted method is called, thus making the code easier to read. We evaluate the effectiveness of these heuristics in Section 4.3.
3.5. Ranking Refactoring Suggestions
The number of the remaining applicable suggestions per host method can be large (on average 7 per method as shown in RQ1) and overwhelm the developer. To prioritize high-quality suggestions, we implemented a ranking mechanism. We use three ranking functions, denoted as $\mathcal{T}_{j}(e_{i})$, where $j$ ranging from 1 to 3. Each function assigns a score to each Extract Method suggestion, resulting in three distinct rankings for each suggestion based on these scores. We present the three ranking functions below:
$\mathcal{T}_{1}(e_{i})$ $=\mathcal{H}(e_{i})=\sum_{i=1}^{N}\mathcal{F}(line_{i})$ $\mathcal{T}_{2}(e_{i})$ $=\mathcal{P}(e_{i})$ $\mathcal{T}_{3}(e_{i})$ $=\mathcal{H}(e_{i})\cdot\mathcal{P}(e_{i})$
@@ 3.5. Ranking Refactoring Suggestions
$N$ $:\text{Total number of lines in the {Extract Method} suggestion.}$.
[…]
$\mathcal{H}(e_{i})$ $:\text{Heat of the {Extract Method} suggestion }e_{i}.$
$\mathcal{P}(e_{i})$ $:\text{Popularity of the {Extract Method} suggestion }e_{i}.$
In Ranking Function 1 ($\mathcal{T}_{1}(e_{i})$), we assign scores to suggestions based on the “heat map” of the host method. Intuitively, this ranking measures the LLM’s confidence that a certain line of code in the host method belongs to another method. We calculate a “heat” score for each applicable refactoring suggestion. To compute the heat map for a specific host method, we follow a systematic process. First, we iterate through each line of the host method, recording how many times each line appears in all applicable suggestions. If a line is absent from all suggestions, we assign it a score of zero. We repeat this procedure for every line within the host method. Subsequently, we aggregate these individual line scores to derive the overall heat score for each suggestion (which comprises several lines of code). Then, we rank suggestions according to their heat score.
Ranking Function 2 ($\mathcal{T}_{2}(e_{i})$) ranks suggestions by taking into account their popularity. Remember that we re-prompt the LLM several times with the same prompt (as discussed in Section 3.1.2). While we remove duplicated suggestions, we keep track of how many times a certain suggestion is produced by the LLM during re-prompting. Given that LLMs have been trained on a substantial volume of source code, enabling them to mimic how real developers construct methods, we give precedence to suggestions that appear repeatedly through this ranking function.
Ranking Function 3 ($\mathcal{T}_{3}(e_{i})$) combines both the heat and popularity of suggestions using a weighted average, aiming to strike a balance between the importance of these two factors. We evaluate the effectiveness of the rankings in Section 4.3.
4. Evaluation
In this section, we empirically evaluate EM-Assist by answering the following research questions:

### v2

@@ III Technique
Figure 2: The workflow of generating refactoring suggestions and then applying them with EM-Assist.
In this section, we present the workflow that our novel approach and tool, EM-Assist, uses to automatically suggest and perform EM refactoring in Java and Kotlin codebases.
Figure 2 shows the architecture and the steps performed by EM-Assist. First, EM-Assist invokes the LLM ( 1 in Figure 2) by prompting it to generate EM suggestions for the selected method (see Section III-A). Next, our algorithm removes hallucinations and retains applicable refactoring options that are free of errors ( 2 in Figure 2, detailed in Section III-B) and are useful ( 3 in Figure 2, detailed in Section III-C). These steps play a key role in retaining suggestions that align with the way developers extract code, while simultaneously mitigating the risk of introducing bugs or non-compilable code.
Next, our algorithm further enhances ( 4 in Figure 2, detailed in Section III-D) the quality of the suggestions by expanding or shrinking the code fragment in the suggestion based on program slicing rules. These rules ensure that refactorings better align with the developer’s intentions. Then, the tool ranks the refactoring suggestions ( 5 in Figure 2, detailed in Section III-E) to prevent overwhelming the developer, presenting only the top-n ranked options for the developer to choose from. Tool’s GUI displays the suggestions and shows a preview of the signature of the extracted method, along with the code fragment that would be extracted. Based on the user’s decision, EM-Assist encapsulates the suggestion in a refactoring command and invokes IntelliJ IDEA to perform the refactoring correctly.
Definition III.1.
(Long method) is chosen by a developer for extract method refactoring, denoted as $l_{i}(s_{i},e_{i})$. The method starts at line $s_{i}$, where its signature is defined, and ends at line $e_{i}$ containing the closing curly bracket that signifies the end of the method declaration. The method’s length, $L_{i}$, is $e_{i}-s_{i}$.
Definition III.2.
@@
(Extract Method) denoted as $e_{j}(n_{j}^{e},(s_{j}^{e},e_{j}^{e})$, where $n_{j}^{e}$ represents the method name, $(s_{j}^{e},e_{j}^{e})$ signifies a code fragment within a long method. It is identified by the starting line number, defined as $s_{j}^{e}>s_{i}$, and the ending line number, defined as $e_{j}^{e}<e_{i}$. These line numbers are calculated with respect to the original source code file in which the host method $l_{i}(s_{i},e_{i})$ is located.
Definition III.3.
@@
(Extract Method Suggestions ($\mathcal{S}$)) is a set of extract method suggestions generated by an LLM for a long method $l_{m}$, formally, $\mathcal{S}=\{e_{i}(n_{i},(s^{e}_{i},e^{e}_{i})),e_{j}(n_{j},(s^{j}_{s},e^{j}_{e})),\ldots\}$
Definition III.4.
@@
(Not Useful Extract Methods $(\mathcal{NU})$) is a subset of $\mathcal{V}_{i}$ comprising elements that fail to meet the criteria for usefulness, $U(e_{i})$. This subset ensures that the suggestions are neither too large, encompassing almost the entire method body, nor too small, essentially one-liners. Formally, $\mathcal{NU}_{i}={e_{i}\in\mathcal{V}\mid\neg U(e_{i})}$. Conversely, useful suggestions $\mathcal{U}_{i}$, consist of elements satisfying the usefulness criteria $U(e_{i})$, suitable for method extraction, defined as $\mathcal{U}_{i}={e_{i}\in\mathcal{V}\mid U(e_{i})}$.
III-A Generating EM Suggestions
@@
We utilize LLMs to generate EM recommendations for a given input long method. In this section, we delve into the preparation of LLM for EM refactoring suggestions, as well as the parameters of the LLM that require tuning.
III-A1 Prompt engineering
@@
LLMs are versatile models capable of various applications, but they require specific preparation when used for a particular task [25]. To facilitate this, we employ in-context learning [33], where we provide all the instructions needed to solve the task right in the model’s input, including task definition and relevant context information. To further enrich the prompt, we employ few-shot learning [32, 34] by incorporating a set of 2 examples into the prompt, following best practices as discussed by Gao et al. [35]. The prompt includes: (i) A succinct overview of EM refactoring, (ii) The definition of a long method, (iii) A JavaDoc string when available, (iv) 2 examples of refactoring for one long and one short method, and (v) Precise instructions for the desired output format. An example prompt is accessible on our companion website [36].
III-A2 Generating an extensive array of suggestions
@@
The output of an LLM depends on two factors: (i) the internal variable “Temperature” ($T$), and (ii) the number of iterations ($I$[32, 37, 34], which indicates how many times the same prompts are inputted. Temperature serves as a regulator for the model’s output randomness. Higher values, such as 0.9 or 1.2, produce more diverse and unpredictable outputs, whereas values closer to 0 produce focused and less diverse results. We determine the optimal value empirically (see Section IV-D).
@@
Even when identical prompts are presented multiple times, the output can exhibit variations due to: (i) LLMs employ a combination of learned patterns and random sampling in generating responses, introducing slight deviations in their answers each time due to inherent randomness; (ii) LLMs possess the capability to explore various potential solutions and refine their responses iteratively through interactions and feedback [32, 34], continually enhancing their outputs in subsequent iterations. Therefore, the quality and quantity of predictions made by an LLM are influenced by the frequency of prompting the same prompt.
In Section IV-D, we conduct a sensitivity analysis to determine how the performance of EM-Assist varies with each combination of temperature and iterations. This analysis enables EM-Assist to use the optimal settings for generating refactoring suggestions.
III-B Removing Invalid EM Suggestions
@@
Not all suggestions generated by LLMs can be directly applied to codebases, as some may lead to non-compilable code. Therefore, EM-Assist analyzes each suggestion to ensure it can be extracted without breaking the code’s scope. For example, suggestion 3 in our motivating example (Figure 1) fails the scope analysis, and EM-Assist expands the code fragment of the suggestion so that both the start and end lines have the same scoping level: the new suggestion’s code fragment starts from line 160, while the end line stays the same. This change, however, leads to a compilation error.
@@
To remove suggestions that might lead to non-compilable code, EM-Assist uses the rules below:
@@
    Variable Usage: This criterion ensures that all necessary variables and objects are accessible within the same scope or are provided as parameters, maintaining the code’s self-containment.
@@
    Return Values: If the code being extracted produces a result or modifies the state of objects, EM-Assist validates that the return value and side effects are appropriately handled. For example, it checks if the return value is used subsequently in the host method or if any thrown exceptions are caught.
  3. (iii)
@@
  4. (iv)
    Control Flow: EM-Assist reasons about the control flow within the extraction suggestion to ensure that it can be extracted as a separate method without causing logical errors. For example, it checks for any early returns or breaks that might affect the control flow and would render the extracted method to have a different behavior than the original code.
EM-Assist checks all these conditions through the static analysis infrastructure in the IntelliJ Platform [^1] The IntelliJ Platform: https://www.jetbrains.com/opensource/idea/. (the open-source platform IntelliJ IDEA and other JetBrains IDEs are built upon) that verifies refactoring preconditions. As LLMs are not aware of the semantics of EM refactoring, but only about the code tokens that are associated together in its training data, this step is crucial for discarding a large number of invalid suggestions (57.4% of all suggestions).
III-C Removing Refactoring Suggestions That Are Not Useful
Once we retain only valid refactoring suggestions, the next step filters out those suggestions that are not useful and retains the ones that are applicable and likely to be performed by a developer. To identify suggestions that are not useful, we employ two rules: (i) Exclude EM suggestions that encompass 88% or more of the statements present in the long method, and (ii) Following practices adopted by other researchers [1, 28, 38, 39], we exclude EM suggestions that are one-liners.
Extracting an entire method or a one-liner into another method is, in general, not useful for the purpose of improving existing code. These might be useful if the developers plan to add new code that implements new features into the host method or into the extracted method. Notice that we designed EM-Assist for code renovation and not for feature expansion.
III-D Enhancing Refactoring Suggestions
After identifying useful suggestions (Definition III.5) devoid of hallucinations, the next step enhances these suggestions to further improve the quality of valid suggestions.
@@ III-D Enhancing Refactoring Suggestions
    Program slicing: Taking inspiration from tools that rely solely on static analysis, such as those utilizing program slicing [7, 1, 8, 40, 41], we leverage program slicing to augment refactoring suggestions provided by LLMs. We use rules that better align with developer preferences to avoid creating small methods with several arguments.
    Let $EF_{s}$ be the initial statement of the code fragment to be extracted into a new method, with $EF_{s+i}$ indicating statements below $EF_{s}$ and $EF_{s-i}$ denoting statements above $EF_{s}$ within the long method. Let $EF_{s-1}$ represent the variable declaration statement for variable $v$, and this variable is used inside the code fragment of the suggested method extraction, identified through live variable analysis based on a def-use chain. In such cases, we increase the scope of the code to be extracted by one statement, including $EF_{s-1}$ in the extracted method, and we adjust the starting line of the suggestion to match the starting line of $EF_{s-1}$. This avoids having to pass $v$ as a parameter.
@@
    Control statements: If the code fragment in the extracted method starts with a control statement (i.e., $EF_{s}$ is the check block of an if statement), we shrink the code fragment to start at $EF_{s+1}$ and contain only the block of the if. This leaves the check condition intact in the host method so that the developer can see under which conditions the would-be extracted method is called, thus making the code easier to read.
We evaluate the effectiveness of these heuristics in Section IV-D. We plan to experiment with other heuristics based on the feedback we receive on JetBrains Marketplace [42].
III-E Ranking Refactoring Suggestions
The number of the remaining applicable suggestions per host method can be large (on average 10 per method) and overwhelm the developer. To prioritize high-quality suggestions, we implemented a ranking mechanism that gives precedence to recommendations consistently identified by the LLM. This approach elevates frequently highlighted suggestions, employing the LLM’s extensive knowledge of coding best practices. Next we present our ranking function.
$\mathcal{T}_{1}(e_{i})$ $=\mathcal{H}(e_{i})=\sum_{i=1}^{N}\mathcal{F}(line_{i})$ (1)
$\mathcal{T}_{2}(e_{i})$ $=\mathcal{P}(e_{i})$ (2)
$\mathcal{T}_{3}(e_{i})$ $=\mathcal{H}(e_{i})\cdot\mathcal{P}(e_{i})$ (3)
@@ III-E Ranking Refactoring Suggestions
$N:$ Total number of lines in the EM suggestion.
$\mathcal{F}(line_{i}):$ Number of times the ith line appears in all the $\text{Number of times the $i^{th}$ line appears in all the suggestions}.$
$\mathcal{H}(e_{i}):$ Heat of the EM suggestion $\text{Heat of the {EM} suggestion }e_{i}.$
$\mathcal{P}(e_{i}):$ Popularity of the EM suggestion $\text{Popularity of the {EM} suggestion }e_{i}.$
Our approach begins with $\mathcal{T}_{1}(e_{i})$, where we assign scores to each suggestion according to a ”heat map” of the host method. Intuitively, this ranking measures the LLM’s confidence that a certain line of code in the host method belongs to another method. To compute the heat map, we record how many times each line appears in all applicable suggestions. If a line is absent from all suggestions, we assign it a score of zero. We repeat this procedure for every line within the host method. Then, we aggregate individual line scores to determine each suggestion’s overall heat score, comprising multiple code lines, and then rank these suggestions based on their heat scores.
$\mathcal{T}_{2}(e_{i})$ ranks suggestions by considering their popularity. As discussed in Section III-A2, we re-prompt the LLM several times with the same prompt. While we remove duplicated suggestions, we keep track of how many times a certain suggestion is produced by the LLM during re-prompting. Given that LLMs have been trained on extensive codebases enabling them to mimic how real developers construct methods, we give precedence to suggestions that appear repeatedly.
Finally, $\mathcal{T}_{3}(e_{i})$ combines both the heat and popularity of suggestions using a weighted average, aiming to strike a balance between the importance of these two factors. We evaluate the effectiveness of the rankings in Section IV-D.
IV Evaluation
We empirically evaluate EM-Assist by answering the following research questions: